OpenStudy (anonymous):

A dump truck can haul 16 loads of dirt in 4 hours. How long will the dump truck spend hauling 25 loads? h

OpenStudy (anonymous):

how much can it haul per hour?

OpenStudy (anonymous):

16/4

OpenStudy (anonymous):

six hours and 15 minutes----16 loads in 4 hours = 4 loads in each hour---25 loads divided by 4 = six and one quarter hours = 375 minutes.

OpenStudy (hanner_b_nanner):

You can use a proportion... 16/4=25/x

OpenStudy (anonymous):

16 / 4 will give you one hour 16 / 4 = 4 Multiply by 25 4 x 25 = 100

OpenStudy (hanner_b_nanner):

but make sure it's turned to minutes now hours :))

OpenStudy (anonymous):

*not